OpenStudy (anonymous):

A rectangle is twice as long as it is wide. If its area is 50 sq. yd., find the length. 10 yd. 25 yd. 33 1/3 yd.

OpenStudy (anonymous):

Let the length be x, wts the breadth?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Alright first start with your area equations The area of a rectangle is \[A=L*W\] where L = Length and W= Width? correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Alright and the first statement says that your length is 2 times the size of the width how can you write this as an equality

OpenStudy (anonymous):

L=? W

OpenStudy (anonymous):

L=2 W

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

yes you're correct

OpenStudy (anonymous):

i am?

OpenStudy (anonymous):

hold on lol

OpenStudy (anonymous):

okay ha

OpenStudy (anonymous):

Right it as W= ? L ... if length is 2x larger than width with is one half of L

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

2W=L or 1/2L=W

OpenStudy (anonymous):

so you were right in the first place i'm just contradicting myself anyways so you have A=L*W L=2W you know the area of the rectangle to be 50 units 50=L*W

OpenStudy (anonymous):

uhh.. i dont get it:(

OpenStudy (anonymous):

You don't get what?

OpenStudy (anonymous):

nevermind.. i get it now

OpenStudy (anonymous):

yeah i just replaced A with 50 since it is given

OpenStudy (anonymous):

next replace L in the area equation 50=2W*W

OpenStudy (anonymous):

\[50=2W^2\] can you solve this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this will get you your width. Then all you need to do is multiply that by 2 for your length

OpenStudy (anonymous):

alright thank you